Gps L3 test 23sep08
There was a gps L3 test on
23sep08. During the test the L3 signal was left on continuously. Time
domain (voltage samples) were taken with the mock spectrometer during
the test. The setup was:
- mock spectrometer run at 172.032Mhz with a decimation of 21
giving a bandwidth of 8.192 Mhz
- I,Q, 12 bits samples, 2 polarization data taken for about 60
- telescope sitting at az=285, za=10.
- alfa beam 0 used. 1379 Mhz was the center of the rf band.
- 16K power spectra were computed and averaged to 40 milliseconds.
800 averaged spectra (32 seconds) were used for the plots/images.
Average spectra (32 seconds) (.ps )
Dynamic spectra of polA (.gif):
- The top plot is polA, the bottom plot is polB
- the freqeuncy resolution is 500 Hz.
- The spectra is the median of the 800 spectra computed.
Dynamic spectra of polB (.gif):
- Each 16K spectra was smoothed and decimated down to 1024 pnts (8
- The median bandpass was divided into each of the 40 msec
spectra to flatten the image. The image shows changes from
this average bandpass.
- The solid horizontal lines in polA spaced every
12 seconds are a radar.